A lightning bolt transfers 22.0 C to the earth. How many electrons are transferred?



1
Expert's answer
2021-01-04T14:34:46-0500

One electron carries the charge of


"e = 1.6\times 10^{-19}C"

Thus, there are


"N = \dfrac{Q}{e}"

electrons in the charge of "Q = 22C". Substituting the values, obtain:


"N = \dfrac{22C}{1.6\times 10^{-19}C} = 1.375\times 10^{20}"

Answer. "1.375\times 10^{20}" electrons.